George A. Madrid 28 November 1995

Critique of “Through the Mirror”

“Through the Mirror” was a fun read. The conflict between the clone and the original is not new, but it is usually presented as a moral or spiritual battle. Is a copy of a spirit as valid as the spirit. I don’t think that I’ve ever seen the conflict on such a purely physical level.

I wanted more details, though. Why is he cloning? Is this common? If so, why does he feel so weird about it? If not, what is driving him to this. What does he do? Couldn’t his occupation relate to his motivations? We don’t see his motivations, only his actions. Why does the idea of the clone having a functioning brain bother him so much. What exactly are the clones used for? Spare parts? or wholesale body replacement? If the latter, wouldn’t it need a functional brain, or do they transplant it?

There are some places where the author seems scared to commit, using weak language which lessens the effect. Two examples from page one, “gulped almost frantically..” (1.3) (If it wasn’t frantic, what was it?) and “he seemed to startle himself.” (1.4) (What did he do that made it seem that way.)

Some things that seemed unlikely or unbelievable to me: The clone-people were going to refund full costs for 20staff that would be required for such a company, their profit-margin must be ridiculously high to cover this, unless they do very high volume, which leads again to some of the questions I asked before. How common is this? Another detail, when Joshua awoke, the orderly didn’t even check on him before leaving the room. This seems wrong, especially since he made noises that could indicate choking.

Does that sort of lobotomy ruin the eye? Won’t they need that later? Again, how are the clones used in this society. What does society think of the clones? Is this legal? If so, what are the laws relating to clones? What if a clone kills its original. What is his legal status.

Oh yes, women don’t have Adam’s apples. (They do have larynges, though.)

The questions raised by the last line were neat. They lead us to wonder how many of these stray clones are wondering around. This would have been more effective if I had known for sure how the clones were used. If the guy in the suit had just had his teeth replaced, well, that not as spectacular. But if it’s a whole clone?